Cull Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on machine-hours. The company based its predetermined overhead rate for the current year on total fixed manufacturing overhead cost of $488,000, variable manufacturing overhead of $2.80 per machine-hour, and 61,000 machine-hours. The company has provided the following data concerning Job X455 which was recently completed:

Number of units in the job 20
Total machine-hours 80
Direct materials $ 880
Direct labor cost $ 1,760

If the company marks up its unit product costs by 20% then the selling price for a unit in Job X455 is closest to: (Round your intermediate calculations to 2 decimal places.)

Garrison 16e Rechecks 2017-06-22

Multiple Choice

  • $175.20

  • $35.04

  • $230.24

  • $210.24
Answer #1

Ans:$210.24

Statement showing Computations

Particulars

Amount

Direct Material per unit= 880/20

                    44.00

Direct Labour per unit= 1760/20

                  88.00

Variable O/H per unit = 2.80*80/20

                    11.20

Fixed Manu o/H per unit = 488,000/61,000*4

                    32.00

Total cost per unit

                  175.20

Mark up = 316.60*.20

                    35.04

selling price for a unit in Job X455

                  210.24